If the earth revolved rapidly enough, the weight of objects at the equator would be zero. What would be the length of the day in that case (in h)?
Answer #1

Circumference of Earth c = 40075.16 km
r = 6371 km
M = 5.98 x 10^24 kg

For m = 0 at center
Force of gravity = Force of Centripetal acceleration
mg = mv^2/r
g = v^2/r
v = sqr(gr)
dx/dt = sqr(gr)
x = c = sqr(gr) x [t] <-- Definite integral between t and 0
t = c/sqr(gr)
t = [40075.16 x 10^3 / sqr(9.8 x 6371 x 10^3)] s.
t = 5071.75 s = 1.408 hours